OpenStudy (anonymous):

(1/27) ^ 3x+2 (simplify)

OpenStudy (anonymous):

\[(\frac{1}{27})^{3x+2}=\frac{1^{3x+2}}{27^{3x+2}}\]\[=\frac{1^{3x}+1^2}{27^{3x+2}}\]\[=\frac{1+1}{27^{3x+2}}\]\[=\frac{2}{27^{3x+2}}\]

OpenStudy (anonymous):

@Zed, the numerator is just 1, not 2, since you wrote \[1^{3x+2}=1^{3x}+1^2\]

OpenStudy (anonymous):

why did u remove the exponents of 1?

OpenStudy (anonymous):

Yes it is, thanks imperialist. My mistake :)

OpenStudy (anonymous):

@shinigami1m: 1 to any power is just 1

OpenStudy (anonymous):

It should be \[\frac{1}{27^{3x+2}}\]

OpenStudy (anonymous):

it can be zero

OpenStudy (anonymous):

No it cannot. Even 1 to the zeroth power is still 1. 1 raised to any complex number (so including the reals) is always 1

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

(1/27)^2x+2 (1^3x*1^2)/27^(3x+2); {1^(3x+2)+1]/[3^(9x+6); (1+1)/[3^(9x+6); final ans: 2/[3^(9x+6)